# Cross-Lingual Consistency of Factual Knowledge in Multilingual Language Models

### v1

@@ Acknowledgements
This publication is part of the project LESSEN with project number NWA.1389.20.183 of the research program NWA-ORC 2020/21 which is (partly) financed by the Dutch Research Council (NWO).
@@
Due to restriction of our GPU resources, we could not test models larger than BLOOM-3b. Extending our analysis to larger-scale models in future work is encouraged to see if the same conclusions reached. Nevertheless, results in Figure 4 indicate that the average CLC grows extremely slowly with the increment of model scale.
@@
Appendix A Prompt-based Probing
@@
Appendix B Ranking-Based Weight
@@
Model CLC Acc en fr nl es ru ja zh ko vi el hu he tr ca ar uk fa
XLM-RoBERTa 32.04 32.54 46.79 37.29 39.41 38.63 26.66 29.42 23.41 31.07 37.07 27.25 31.33 26.31 37.03 33.19 29.73 28.74 29.81
mT5-large 32.31 33.37 47.44 36.84 40.46 41.64 33.38 22.22 18.9 32.02 34.07 24.84 30.26 29.46 41.78 36.56 34.63 32.92 29.95
LLaMA-7b 28.52 28.51 51.69 31.33 41.48 38.55 34.54 19.92 15.15 19.04 22.29 20.38 28.43 23.47 23.22 34.31 23.07 35.51 22.29
BLOOM-7.1b 26.26 27.82 56.39 35.72 27.03 45.32 22.25 20.35 28.62 18.15 41.92 20.35 16.84 23.29 21.14 33.11 24.68 22.63 15.22
BLOOMz-3b 24.32 23.87 52.34 37.53 20.57 39.81 16.46 18.62 23.95 17.21 37.57 15.53 16.87 16.11 16.48 24.01 21.32 18.20 13.22
BLOOM-3b 25.18 25.97 54.74 34.57 23.35 40.95 22.10 20.88 25.40 19.36 37.66 20.07 16.18 20.23 20.32 28.30 21.42 22.67 13.31
BLOOM-1.7b 24.83 24.77 51.53 31.83 22.73 38.68 24.01 20.14 23.31 17.02 35.92 20.27 16.20 19.26 19.98 25.07 19.86 20.36 14.84
BLOOM-1.1b 23.93 22.93 48.07 28.00 20.89 36.45 22.61 18.91 21.55 16.08 31.30 19.30 16.14 17.73 17.08 22.20 17.80 21.91 13.77
BLOOM-560m 23.10 21.14 42.15 25.90 18.11 29.67 23.12 18.14 19.11 16.62 26.74 19.05 16.64 19.41 16.02 19.60 14.93 21.01 13.19
Table 4: Overall PLM probing accuracy. Probing accuracy raises with the increment of model size in BLOOM series.
Appendix C RankC Computation Example
@@
Figure 8: RankC scores between language pairs in BMLAMA-53.
Figure 10: Genetic similarity between language pairs in BMLAMA-53.
Appendix D Range of RankC Values
In practice, if a model outputs the same answers (the ones with the highest probabilities) for all queries in two languages, the $P@1$ values for each query are constantly 1 and the RankC value between these two languages should be equal or larger than $\frac{\sum_{i=1}^{|Q_{l}|}w_{1}*P@1}{|Q_{l}|}=\frac{\sum_{i=1}^{|Q_{l}|}w_{1}}{|Q_{l}|}$, where $w_{1}$ is the weight for the candidates with the highest probability, which varies along with the length of the candidate sets. However, since the average candidate is about 10 for BMLAMA, $w_{1}$ can be approximated as $\frac{e^{9}}{e^{0}+\ldots+e^{9}}$ for nearly all queries. The RankC score under this ‘all-same-answer’ hypothesis is
Therefore, two languages are considered as higher CLC if their RankC scores are closer to or larger than 63%.
Appendix E Comparison of RankC and COverlap
@@
Appendix F Effect of Supported Language in BLOOM
Despite the official supported languages of BLOOM, we decided to include all languages in BMLAMA because we found the factual knowledge probing accuracy of the non-supported languages to be considerably higher than randomly guessing (cf. Table  4). As suggested by previous study (Blevins and Zettlemoyer, 2022), this could be due to the actual presence in the training data of some non-officially supported languages.
To enhance the convincement of our experiment on BLOOM, we further calculate the correlation scores between CLC and word-overlapping when only considering the supported languages.
Shown in Table 5, the correlation scores between CLC and sub-word vocabulary overlap become even higher, confirming our finding that such overlap is a key predictor of consistency across languages.
@@ Appendix F Effect of Supported Language in BLOOM
Appendix G Additional Results on BMLAMA-53
@@
The genetic similarity in Figure 10 gives a brief view of connections with CLC. For instance, the Russian-Ukrainian pair achieve 80% genetic similarity while it also shows high CLC in all PLMs. Figure 9 shows the subword vocabulary overlap scores on BMLAMA-53. We observe that high subword vocabulary overlapping leads to high RankC scores, such as sr-uk/be-uk/bg-uk/bg-sr/sk-sc pairs in BLOOM-3b, sl-hr/sv-da/id-ms pairs in mT5-large, and gl-es/id-ms/pt-gl pairs in XLM-RoBERTa-large.
@@
Appendix H Extra Results on More Models
To enhance the comprehension of our study, we conducted additional experiments on more models: BLOOM-7.1b, a larger model in BLOOM series; BLOOMz-3b, a branch fine-tuned with human instruction; LLaMA-7b, one of the cutting-edge PLMs. For replicability, we provide the values of average CLC and probing accuracy for all PLMs in Table 4.
For BLOOM-7.1b, its probing accuracy raises 1.85%, while the average CLC still grows relatively slowly (1.08%) in BLOOM series. For BLOOMz-3b, both the average CLC and probing accuracy are lower than BLOOM-3b, suggesting human instruction fine-tuning has a mildly negative effect on factual knowledge. Nevertheless, average cross-lingual consistency (CLC) still remains less affected, and the heatmaps are nearly the same as BLOOM-3b since they use the exact same tokenizer. This further strengthens our conclusion that CLC is a property independent of accuracy and mostly relies on the subword overlapping. For the probing accuracy of LLaMA-7b, it only slightly outperforms BLOOM-7.1b, which has a similar number of parameters. For average CLC, LLaMA-7b reaches 28.52% and BLOOM-7.1b reaches 26.26%. In addition, we calculate the Pearson correlation scores between CLC and vocabulary overlapping of these two models. Results in Table 6 also support our finding that CLC has a positive correlation with the overlapping of sub-word vocabulary.
@@ Appendix H Extra Results on More Models
Appendix I Additional Cases of Counterfactual Knowledge Incorporation
@@
Three extra queries are shown in Table 7. The languages with high RankC scores with English (i.e. es, vi) also change synchronously when incorporating new knowledge into BLOOM-3b. Meanwhile the languages with low RankC scores are less affected, in line with our findings in Section 7.
@@
Appendix J Raw Logits Before/After Model Editing
For replicability, Table J shows the raw (unnormalized) logits of the model editing case study.

### v5

@@ Acknowledgements
This publication is part of the project LESSEN with project number NWA.1389.20.183 of the research program NWA-ORC 2020/21 which is (partly) financed by the Dutch Research Council (NWO). We also gratefully acknowledge the computational resources provided by the High-Performance Computing cluster Hábrók. RF is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 819455).
@@
Due to restriction of our GPU resources, we could not test models larger than BLOOM-7.1b. Extending our analysis to larger-scale models in future work is encouraged to see if the same conclusions reached. Nevertheless, results in Figure 4 indicate that the average CLC grows extremely slowly with the increment of model scale.
@@
Appendix A GenBench Evaluation Card
Figure 4 shows the evaluation card for our experiments on BMLAMA according to the generalization taxonomy of Hupkes et al. (2022).
Motivation
Practical Cognitive Intrinsic Fairness $\square$ $\square$
Generalisation type
Compo- sitional Structural Cross Task Cross Language Cross Domain Robust- ness $\square$
Shift type
Covariate Label Full Assumed $\square$
Shift source
Naturally occuring Partitioned natural Generated shift Fully generated $\square$
Shift locus
Train–test Finetune train–test Pretrain–train Pretrain–test $\square$
Table 4: Evaluation card for experiments on BMLAMA.
Appendix B Prompt-based Probing
@@
Appendix C Ranking-Based Weight
@@
Appendix D RankC Computation Example
@@
Appendix E Range of RankC Values
In practice, if a model outputs the same answers (the ones with the highest probabilities) for all queries in two languages, the $P@1$ values for each query are constantly 1 and the RankC value between these two languages should be equal or larger than $\frac{\sum_{i=1}^{|Q_{l}|}w_{1}*P@1}{|Q_{l}|}=\frac{\sum_{i=1}^{|Q_{l}|}w_{1}}{|Q_{l}|}$, where $w_{1}$ is the weight for the candidates with the highest probability, which varies with the length of the candidate sets. However, since the average candidate is about 10 for BMLAMA, $w_{1}$ can be approximated by $\frac{e^{9}}{e^{0}+\ldots+e^{9}}$ for nearly all queries. The RankC score under this ‘all-same-answer’ hypothesis is then
Therefore, two languages are considered highly consistent if their average RankC score is around 63% or larger.
@@ Appendix E Range of RankC Values
Appendix F Comparison of RankC and COverlap
@@
Model CLC Acc en fr nl es ru ja zh ko vi el hu he tr ca ar uk fa
RoBERTa 32.0 32.5 46.8 37.3 39.4 38.6 26.7 29.4 23.4 31.1 37.1 27.3 31.3 26.3 37.0 33.2 29.7 28.7 29.8
mT5-large 32.3 33.4 47.4 36.8 40.5 41.6 33.4 22.2 18.9 32.0 34.1 24.8 30.3 29.5 41.8 36.6 34.6 32.9 30.0
LLaMA-7b 28.5 28.5 51.7 31.3 41.5 38.6 34.5 19.9 15.2 19.0 22.3 20.4 28.4 23.5 23.2 34.3 23.1 35.5 22.3
BLOOM-7.1b 26.3 27.8 56.4 35.7 27.0 45.3 22.3 20.4 28.6 18.2 41.9 20.4 16.8 23.3 21.1 33.1 24.7 22.6 15.2
BLOOMZ-3b 24.3 23.9 52.3 37.5 20.6 39.8 16.5 18.6 24.0 17.2 37.6 15.5 16.9 16.1 16.5 24.0 21.3 18.2 13.2
BLOOM-3b 25.2 26.0 54.7 34.6 23.4 41.0 22.1 20.9 25.4 19.4 37.7 20.1 16.2 20.2 20.3 28.3 21.4 22.7 13.3
BLOOM-1.7b 24.8 24.8 51.5 31.8 22.7 38.7 24.0 20.1 23.3 17.0 35.9 20.3 16.2 19.3 20.0 25.1 19.9 20.4 14.8
BLOOM-1.1b 23.9 22.9 48.1 28.0 20.9 36.5 22.6 18.9 21.6 16.1 31.3 19.3 16.1 17.7 17.1 22.2 17.8 21.9 13.8
BLOOM-560m 23.1 21.1 42.2 25.9 18.1 29.7 23.1 18.1 19.1 16.6 26.7 19.1 16.6 19.4 16.0 19.6 14.9 21.0 13.2
Table 5: Overall probing results. Probing accuracy rises with the increment of model size in BLOOM series.
Appendix G Effect of Non-Supported Languages in BLOOM
BLOOM does not officially support all languages in BMLAMA. Despite that, we found the factual knowledge probing accuracy of the non-supported languages to be considerably higher than randomly guessing (cf. Table  5) and thus decided to include all BMLAMA languages in our main CLC evaluation. As suggested by previous work (Blevins and Zettlemoyer, 2022), this could be due to the actual presence in the training data of some non-officially supported languages.
To ensure the presence of non-supported languages did not affect our analysis of BLOOM, we further calculate the correlation scores between CLC and word-overlapping when only considering the supported languages. As shown in Table 6, the correlation scores between CLC and sub-word vocabulary overlap become even higher, confirming our finding that such overlap is a key predictor of consistency across languages.
@@ Appendix G Effect of Non-Supported Languages in BLOOM
Figure 8: RankC scores between language pairs in BMLAMA-53.
Figure 10: Genetic similarity between language pairs in BMLAMA-53, calculated by lang2vec (Littell et al., 2017) based on the distance of two languages in the Glottolog tree (Hammarström et al., 2017).
Appendix H Additional Results on BMLAMA-53
@@
Figure 10 shows the pairwise genetic similarity scores in BMLAMA-53. For instance, the Russian-Ukrainian pair here achieves 80% genetic similarity while it also shows high CLC in all PLMs.
Figure 9 shows the subword vocabulary overlap scores in BMLAMA-53. We observe that high subword vocabulary overlapping leads to high RankC scores, such as sr-uk/be-uk/bg-uk/bg-sr/sk-sc pairs in BLOOM-3b, sl-hr/sv-da/id-ms pairs in mT5-large, and gl-es/id-ms/pt-gl pairs in XLM-RoBERTa-large.
@@
Appendix I Knowledge Consistency on More Models
We conducted additional experiments on more models than those included in the main body: namely, BLOOM-7.1b, a larger model in the BLOOM series; BLOOMZ-3b (Muennighoff et al., 2023), a branch of BLOOM-3b fine-tuned with human instruction; LLaMA-7b (Touvron et al., 2023), one of the cutting-edge PLMs trained on much more tokens (i.e. 1.4T) than BLOOM (i.e. 366B). Average CLC and probing accuracy for all PLMs are presented in Table 5.
BLOOM-7.1b: Probing accuracy raises 1.8% relative to BLOOM-3b, while the average CLC still grows relatively slowly (1.1%) in the BLOOM series.
BLOOMZ-3b: both average CLC and probing accuracy are lower than BLOOM-3b, suggesting human instruction fine-tuning has a mildly negative effect on factual knowledge. Nevertheless, average cross-lingual consistency (CLC) still remains less affected, and the heatmaps are nearly the same as BLOOM-3b since they use the exact same tokenizer. This further strengthens our conclusion that CLC is a property independent of accuracy and mostly relies on the subword overlapping.
LLaMA-7b: Probing accuracy only slightly outperforms BLOOM-7.1b having a similar number of parameters. For average CLC, LLaMA-7b reaches 28.5% while BLOOM-7.1b reaches 26.3%. In addition, we calculate the Pearson correlation scores between CLC and vocabulary overlapping of these two models. Results in Table 7 also support our finding that CLC positively correlates with the overlapping of sub-word vocabulary.
@@ Appendix I Knowledge Consistency on More Models
Appendix J Additional Cases of Counterfactual Knowledge Incorporation
@@
Three extra queries are shown in Table 8. The languages with high RankC scores with English (i.e. es, vi) also change synchronously when incorporating new knowledge into BLOOM-3b. Meanwhile, the languages with low RankC scores are less affected, in line with our findings in Section 7.
@@
Appendix K Raw Logits Before/After Model Editing
For replicability, Table K shows the raw (unnormalized) logits of the model editing case study.